OpenStudy (anonymous):

integrating exponential functions x^4e^(x^5) dx

OpenStudy (anonymous):

i kno wu=x^5

OpenStudy (anonymous):

\[\int\limits x^4e^{x^5} dx\] Let u =\[x^{5}\] So, the integral is now: \[\int\limits x ^{4} e ^{u} dx\] u=\[x^{5}\] so du=\[5x^{4} dx\] and dx=du/(5x^4) Integral: \[\int\limits x^{4} e^{u} dx/(5x^{4})\] The x^4 factors cancel out, and you can pull the (1/5) outside of the integral. All you're left with now is \[1/5 \int\limits e^{u} du = (1/5)e^{u} + C\] Plugging back in for u gives the answer as: \[(1/5)e^{x^5}+C\]

OpenStudy (anonymous):

x^4e^(x^5) dx u = x^5 --> du/5 = x^4dx = (1/5) Int e^u du = (1/5) e^(x^5) + C